Use graphing calculator to explain answer please. Round results to the nearest 4 decimal places(3 points) A news report states that the 99% confidence interval for the mean number of daily calories consumed by participan
Answer #1

Answer:

Given,

99% confidence interval for population mean = (xbar - s/sqrt(n)*t (0.005 , n-1) , xbar + s/sqrt(n)*t (0.005 , n-1)

n = 15

xbar - s/sqrt(15)*t (0.005 , 14) = 1870

xbar + s/sqrt(15)*t (0.005 , 14) = 2120

So,

a)

Sample mean = (1870 + 2120)/2

= 1995

b)

Margin of error = (2120 - 1870)/2

= 125

c)

So consider,

Margin of error = s/sqrt(15)*t(0.005,14) = 125

s/sqrt(15)*2.977 = 125

s = 125*sqrt(15) / 2.977

s = 162.62